two angles are complementary. their difference is 36 degrees. find the angles.
Answer by stanbon(75887)

two angles are complementary. their difference is 36 degrees. find the angles.
----
Equations:
x + y = 90
x - y = 36
--------
Add and solve for "x":
2x = 116
x = 58
---
Solve for "y":
x+y = 90
58 + y = 90
y = 32
-----------------
Cheers,
Stan H.